OpenStudy (anonymous):

George has found a relationship between the height of a person and his bowling score. The table below shows the data collected by George: Height cm 152 160 164 170 180 162 154 158 168 Bowling Score 51 55 57 60 65 55 52 54 59 Part A: What would most likely be the bowling score of a person who has a height of 156 cm? (3 points) Part B: Predict a possible correlation coefficient for the data in the table and explain why you think your prediction is a good value for the data. (4 points) i Just need help with B nothing else

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

Well you can plot the data and estimate a value for it. A strong correlation would be probably about 0.8-1 A low correlation maybe 0-0.4 It doesn't seem like they care from the question for an exact value, so I think you could just eye-ball it

OpenStudy (kirbykirby):

Otherwise software would be your best bet to get the correlation coefficient.

OpenStudy (anonymous):

what do you mean software?

OpenStudy (dangerousjesse):

Bowling score increases with height. Try a linear regression fit. Since 156 is between 154 and 158, try looking between the corresponding bowling scores.

OpenStudy (dangerousjesse):

But you're going to have to put the heights and scores in order from least to greatest.

OpenStudy (anonymous):

ok but i just learned what a correlation coefficient is and i still dont understand it that well, can someone explain it?

OpenStudy (anonymous):

and i have no idea how to do a linear regression fit

OpenStudy (dangerousjesse):

It's pretty much just a wordy way of saying "Look at how the variables fit in with each other when the numbers are in order."

OpenStudy (anonymous):

ok and what about the other one

OpenStudy (kirbykirby):

 On the left: the fitted line (a.k.a. regression line) fits through the poiints much better than on the right, so the left diagram would have a higher correlation coefficient than the right one

OpenStudy (anonymous):

ok but im still confused on how to find the correlation coefficient

OpenStudy (kirbykirby):

Hopefully this will give more useful explanation: http://onlinestatbook.com/2/describing_bivariate_data/pearson.html

OpenStudy (dangerousjesse):

The correlation coefficient, aka the cross-correlation coefficient, is a quantity that gives the quality of a least squares fitting to the original data. To really explain the correlation coefficient, you need to consider the sum of squared values \[ss_{xx}\]\[ss_{xy}\] and \[ss_{yy}\] of a set of n data points \[(x_{i}, y{i})\] about their respective means, \[ss_{xx}\]\[= \sum (x_{i}-x)^{2}\]\[= \sum x^{2} - 2 x \sum x + \sum x^{2}\]\[=\sum x^{2} -2 x \sum x + \sum x^{2}\]\[=\sum x^{2}-nx^{2}\] and so on and so forth. These quantities are basically unnormalized forms of the variances and covariance of X and Y given by \[ss_{xx} = N var (x)\]\[ss_{yy} = N var (Y)\]\[ss_{xy} = N cov (x,y)\]

OpenStudy (dangerousjesse):

Sorry for taking so long, I hate writing equations on hand -.-

OpenStudy (anonymous):

ok i think i got it can you help me with one more thing

OpenStudy (dangerousjesse):

Sure